A sum of money is divided between Raymond and Roland in the ratio of 3:5. Roland gets $300 more than Raymond. What is the sum of money divided between them?
Answer by ankor@dixie-net.com(22740)

A sum of money is divided between Raymond and Roland in the ratio of 3:5.
Let x = the multiplier, then
3x = Ray's $
5x = Rol's $
" Roland gets $300 more than Raymond."
5x = 3x + 300
5x - 3x = 300
2x = 300
x = 300/2
x = 150 is the multiplier, therefore
3(150) = $450 for Ray
and
5(150) = $750 for Rol
:
What is the sum of money divided between them?
450 + 750 = $1200